HI HERES MY PROBLEM
My husband Is 4 times older than my daughter now. it 20 years he'll be twice as old.
how old is the father?

Answer by Theo(2973)

h = husband's age
d = daughter's age.

Your husband is 4 times your daughter's age now>

This is represented by the formula:

h = 4d

Your husband will be 2 times as old as your daughter in 20 years.

This is represented by the formula:

(h+20) = 2(d+20)

Since h = 4d from the first equation, then substitute 4d for h in the second equation to get:

(4d + 20) = 2(d+20)

you now have one equation in one unknown that you can solve.

simplify to get:

4d + 20 = 2d + 40

subtract 2d from both sides of the equation and subtract 20 from both sides of this equation to get:

2d = 20

divide both sides of the equation by 2 to get:

d = 10

Your daughter is 10 years old now.

Since h = 4d, then your husband is 40 years old now.

In 20 years your daughter will be 30

In 20 years your husband will be 60.

In 20 years your husband will be twice your daughter's age.

Your answer is that the father is 40 years old now.